# Cluster Magnification, Root Capacity, Unique Chains and Base Change

### v2

@@ Abstract.
This article is inspired from the work of Krithika and Vanchinathan in [5] and the work of Alexander Perlis in [7] and [8]. We establish the existence of polynomials for given degree and cluster size over number fields which generalises a result of Perlis. We state the Strong cluster magnification problem and establish an equivalent criterion for that. We provide an important example answering a question about Cluster Towers. We introduce the concept of Root capacity and prove some of its properties. We also introduce the concept of unique descending and ascending chains for extensions and establish some properties and explicitly compute some interesting examples. Finally we establish results about all these phenomena under a particular type of base change. The article concludes with results about strong cluster magnification and unique chains.
@@ 1. Introduction
In Sec. 3, we generalise a result of Perlis for number fields that also improves on the generalisation proved previously by Krithika and Vanchinathan in [5]. This generalisation, Thm. 3.1.1 establishes existence of polynomials for given degree and cluster size over number fields. In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very applicable in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
In Sec. 4, we state the Strong cluster magnification problem and establish the following equivalent criterion for that in Thm. 4.1.5 in terms of Galois groups. For all notations, see Sec. 4.
@@
In Sec. 6, we introduce the concept of Root capacity. We begin the section with some observations about the group of automorphisms of finite extensions in Prop 6.1.1 and Prop 6.1.3 and its corollaries. Then we prove some properties of root capacity in propositions 6.2.2, 6.2.4, 6.2.6, 6.2.8 and 6.2.9. The section concludes with Thm 6.2.10.
In Sec. 7, we introduce the concept of unique descending and ascending chains for extensions. Thm. 7.1.1 and Thm. 7.2.1 encapsulate the important properties of unique chains. We compute unique descending chains for some interesting examples Example 7.1.4, Example 7.1.6 and Thm. 7.1.7. Finally Prop. 7.2.2 links unique ascending and descending chain under certain conditions.
In Sec. 8, we establish results about cluster magnification, root capacity and unique chains under a particular type of base change. Thm. 8.1.1 is base change theorem for strong cluster magnification. Thm. 8.1.8 is base change theorem for weak cluster magnification. Thm. 8.2.2 is about root capacity under base change. Thm. 8.3.1 is about unique chains under base change. The article concludes with Thm. 8.4.1 and Thm. 8.4.2 about strong cluster magnification and unique chains.
@@ 2.1. Root clusters
    $r_{K}(f)=$ number of roots fixed by $H=|{\rm Aut}(K(\alpha)/K)|=[N_{G}(H):H]$.
@@ 3.1. Existence of polynomials for given degree and cluster size over number fields
Reader is referred to Völklein [9] for details. Let $K$ be a number field. The field $\mathbb{Q}$ is hilbertian. ([9, Def. 1.9 and Theorem 1.23]). From [9, Corollary 1.11 ] we conclude that every finitely generated extension of a hilbertian field is hilbertian. Hence $K$ is hilbertian. We know that the group ${\mathfrak{S}}_{n}$ occurs regularly over every field ([9, Def. 1.14 and Example 1.17]). In particular, ${\mathfrak{S}}_{n}$ occurs regularly over $K$. The result [9, Corollary 1.15] says that if a group occurs regularly over a hilbertian field then it is realisable as a Galois group over that field. Hence finally we conclude that ${\mathfrak{S}}_{n}$ is realisable as Galois group over $K$.
The following lemma is the final proposition in Perlis [7]. We write Perlis’ proof for the sake of completion.
@@
Let $G$ act transitively on $n$ symbols $\{1,2,\ldots,n\}$. Let $H\subset G$ be the stabiliser of the symbol $1$. Then there is a canonical labelling of the $n$ cosets in $G/H$ so that $G$ acts on $G/H$ exactly the same way $G$ acts on the original $n$ symbols that is $\{x_{1}H,x_{2}H,\ldots,x_{n}H\}$ with $x_{j}\cdot 1=j$ and $g\cdot(x_{j}H)=x_{g\cdot j}H$ for all $g\in G$ and all $1\leq j\leq n$. The action is faithful and transitive. We have $Stab(x_{j}H)=x_{j}Hx_{j}^{-1}$. Hence, $\bigcap\limits_{1\leq j\leq n}(x_{j}Hx_{j}^{-1})=\bigcap\limits_{1\leq j\leq n}({\rm Stab}(x_{j}H))$ is trivial since action is faithful.
Let $N/K$ be finite Galois extension with Galois group isomorphic to $G$. We identify it with $G$. Let $L$ be the subfield of $N$ fixed by $H$. Since the Galois closure of $L/K$ in $N$ is the subfield of $N$ corresponding to the intersection of the conjugates of $H$ in $G$ and that intersection is trivial because of argument in previous paragraph. Hence $N$ is the Galois closure of $L/K$. Let $f$ over $K$ be the minimal (hence irreducible) polynomial for a primitive element of $L/K$. We can identify the $n$ roots of $f$ with the $n$ cosets in $G/H$. This is the required $f$. ∎
@@
As a corollary, we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$. Corollary also holds for $G^{n}$ in place of $G$ for any $n\in\mathbb{N}$. Corollary also holds for $K^{\prime}$.
@@
Now for $i\neq j$, $E_{i}\cap E_{j}$ corresponds to subgroup generated by $N_{i}$ and $N_{j}$ which is $G^{n}$. Hence $E_{i}\cap E_{j}=K$. Suppose for some $i\neq j$ we have $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}$. Since $E_{i}\cap E_{j}=K$. Hence $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}=K$. Now suppose that all $E_{i}\cap K^{\prime}$ are distinct. Since we have $n$ intermediate fields of $K^{\prime}/K$, $E_{i}\cap K^{\prime}=K$ for some $i$. In either case we get an $i$ such that $E_{i}\cap K^{\prime}=K$. Hence ${\rm Gal}(E_{i}K^{\prime}/K^{\prime})\cong{\rm Gal}(E_{i}/K)\cong G$. This realises $G$ as a Galois group over $K^{\prime}$. By replacing $G$ with $G^{m}$ for any $m\in\mathbb{N}$ in the above argument, we can realise $G^{m}$ over $K^{\prime}$ for any $m$.
For the proof of the corollary, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$. Since $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$. Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$ with $G$ as Galois group of each $E_{i}/K$. ∎
@@
Now suppose $r>1$. In solutions of Exercises 3 and 4 in [7], a solvable group $G\subseteq{\mathfrak{S}}_{n}$ is constructed with the properties that its action is transitive on $n$ points, and a point stabiliser fixes precisely $r$ points. The construction is as follows: We divide the $n$ points into $n/r=s$ packets of size $r$. Let $G$ be the group of permutations on these $n$ points generated by independent cyclic permutations on each packet, together with a cyclic permutation on the overall set of packets. Hence $G$ is transitive. This construction of $G$ has the explicit description of a semidirect product of an $s$-fold direct product of cyclic groups ${\mathbb{Z}}/r{\mathbb{Z}}$ and a cyclic group ${\mathbb{Z}}/s{\mathbb{Z}}$.
$G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}},$
where $\text{where each}\ {\mathbb{Z}}/r{\mathbb{Z}}\ \text{permutes points in a packet and action of}\ {\mathbb{Z}}/s{\mathbb{Z}}\ \text{permutes}\ s\ \text{copies of}\ {\mathbb{Z}}/r{\mathbb{Z}}.$
@@
Since direct product of solvable groups is solvable, direct products $G^{i}$ for $i\in\mathbb{N}$ are solvable. By Shafarevich’s theorem, $G^{i}$ for $i\in\mathbb{N}$ are realisable as Galois groups over $\mathbb{Q}$. Hence by Lemma 3.1.4, $G$ is realisable as Galois group over number field $K$. By Lemma 3.1.3, there exists an irreducible polynomial $f$ over $K$ of degree $n$ and a labelling of the roots so that the Galois group of $f$, viewed as a group permuting roots of $f$, is precisely $G$. This $f$ satisfies $r_{K}(f)=r$.
@@ 3.2. A simple lemma about $s_{K}(L)$
We begin this subsection by giving an alternate proof for last equality in (3) of Thm 2.1.1, which is stated in [8] and proved in first Proposition in unpublished note of Perlis [7]. The equality states: $r_{K}(f)=[N_{G}(H):H].$
@@
Let $L_{1},L_{2},\dots,L_{s_{K}(L)}$ be the distinct fields as in above lemma. Hence we have $\tilde{L}=L_{1}L_{2}\dots L_{s_{K}(L)}$, that is the Galois closure of $L/K$ is compositum of distinct fields isomorphic to $L$ over $K$.
@@
Let $K$ be a number field. Note that ${}^{n}P_{k}$ and ${}^{n}C_{k}$ are integers with ${}^{n}P_{k}=k!$ ${}^{n}C_{k}$. By Thm. 3.1.1, we get irreducible polynomial over the field with degree ${}^{n}P_{k}$ and cluster size $k!$. The following theorem says that this is true under some condition also for a general perfect field.
By using the above lemma we give an alternate proof for Thm 3 in [5].
@@
Let $K$ be a perfect field. Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$ (For $K$ number field such $f$ always exists. See $r=1$ case in Proof of Thm 3.1.1).
For $k<n-1$, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$. Let $g$ be the irreducible polynomial over $K$ for a primitive element of $L_{k}$. This polynomial has degree ${}^{n}P_{k}$ and cluster size $k!$.
@@
Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$ be roots of $f$ in $\bar{K}$. Let $L_{k}=K(\alpha_{1},\alpha_{2},\ldots,\alpha_{k})$. We have that degree of $L_{k}/K$ is $n(n-1)\ldots(n-k+1)=$${}^{n}P_{k}$. Since Galois group of $f$ is ${\mathfrak{S}}_{n}$, we have $L^{\prime}$ is isomorphic to $L_{k}$ over $K$ $\iff$ $L^{\prime}=K(\alpha_{i_{1}},\alpha_{i_{2}},\ldots,\alpha_{i_{k}})$ for $k$ roots $\alpha_{i_{j}}$ of $f$. By Lemma 3.2.1, $s_{K}(L_{k})$ is number of distinct fields inside $\bar{K}$ isomorphic to $L_{k}$ over $K$ which is precisely the number of ways of choosing $k$ roots from $n$ roots which is ${}^{n}C_{k}$. By Thm 2.1.1 (2), $r_{K}(L_{k})=k!$. ∎

### v3

@@ Abstract.
This article is inspired from the work of M Krithika and P Vanchinathan in [5] and the work of Alexander Perlis in [7] and [8]. We establish the existence of polynomials for given degree and cluster size over number fields which generalises a result of Perlis. We state the Strong cluster magnification problem and establish an equivalent criterion for that. We also discuss the notion of weak cluster magnification and prove some properties. We provide an important example answering a question about Cluster Towers. We introduce the concept of Root capacity and prove some of its properties. We also introduce the concept of unique descending and ascending chains for extensions and establish some properties and explicitly compute some interesting examples. Finally we establish results about all these phenomena under a particular type of base change. The article concludes with results about strong cluster magnification, unique chains and some properties of the ascending index for a field extension.
@@ 1. Introduction
In Sec. 2, we set up some notations and review the results by Perlis in [7], and Krithika-Vanchinathan in [8], [5].
In Sec. 3, we generalise a result of Perlis for number fields that also improves on the generalisation proved previously by Krithika and Vanchinathan in [5]. This generalisation, Thm. 3.1.1 establishes existence of polynomials for given degree and cluster size over number fields. In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very useful in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
In Sec. 4, we state the Strong cluster magnification problem and establish the following equivalent criterion for that in Thm. 4.1.6 in terms of Galois groups. For all notations, see Sec. 4.
@@
In Sec. 6, we introduce the concept of Root capacity. We begin the section with some observations about the group of automorphisms of finite extensions in Prop. 6.1.1 and Prop. 6.1.3 and its corollaries. Then we prove some properties of root capacity and conclude with Thm. 6.2.13.
In Sec. 7, we introduce the concept of unique descending and ascending chains for extensions. Thm. 7.1.1 and Thm. 7.2.1 encapsulate the important properties of unique chains. Prop. 7.2.2 links unique ascending and descending chain under certain conditions. We prove an interesting result Prop. 7.1.4 that describes the field $N_{1}$ in unique descending chain in terms of the sums of elementary symmetric functions. In Sec. 7.3, we compute unique ascending/ descending chains for some interesting examples (see Prop. 7.3.1, Thm. 7.3.4 and Thm. 7.3.6).
In Sec. 8, we establish results Thm. 8.1.1, 8.1.9, 8.2.2 and 8.3.1 about strong/weak cluster magnification, root capacity and unique chains, respectively under a particular type of base change. We prove results Thm. 8.4.1 and Thm. 8.4.3 about strong cluster magnification and unique chains.
The article concludes with Sec. 9 with some properties of the ascending index $t_{K}(L)$ defined in Thm. 7.2.1 in the context of unique ascending chain for an extension $L/K$. The ascending index has many properties similar to the cluster size $r_{K}(L)$ but has no immediate description in terms of roots of the minimal polynomial of $\alpha$ over $K$ when $L=K(\alpha)$. In Sec. 9, Prop. 9.0.3 establishes a base change property for $t_{K}(L)$, Prop. 9.0.4 establishes an analogue of Cluster Magnification Theorem (Thm. 2.2.3) for $t_{K}(L)$ and Thm. 9.0.5 is an analogue of Thm. 3.1.1 and establishes existence of an extension $L/K$ with prescribed degree and $t_{K}(L)$.
@@ 2.1. Root clusters
    $r_{K}(f)=$ number of roots of $f$ fixed by $H=|{\rm Aut}(K(\alpha)/K)|=[N_{G}(H):H]$.
@@ 3.1. Existence of polynomials for given degree and cluster size over number fields
The reader is referred to Völklein [10] for details. The field $\mathbb{Q}$ is hilbertian (see [10, Def. 1.9 and Theorem 1.23]). Furthermore, every finitely generated extension of a hilbertian field is hilbertian (see [10, Corollary 1.11 ]), and thus we conclude that every number field is hilbertian.
Let $K$ be a number field. We know that the group ${\mathfrak{S}}_{n}$ occurs regularly over every field ([10, Def. 1.14 and Example 1.17]). In particular, ${\mathfrak{S}}_{n}$ occurs regularly over $K$. The result [10, Corollary 1.15] says that if a group occurs regularly over a hilbertian field, then it is realisable as a Galois group over that field. Hence finally we conclude that ${\mathfrak{S}}_{n}$ is realisable as Galois group over $K$.
The following lemma is the final proposition in Perlis [7]. We write the proof given by Perlis for the sake of completion.
@@
Let $G$ act transitively on $n$ symbols $\{1,2,\ldots,n\}$. Let $H\subset G$ be the stabiliser of the symbol $1$. Then there is a canonical labelling of the $n$ cosets in $G/H$ so that $G$ acts on $G/H$ exactly the same way $G$ acts on the original $n$ symbols that is $\{x_{1}H,x_{2}H,\ldots,x_{n}H\}$ with $x_{j}\cdot 1=j$ and $g\cdot(x_{j}H)=x_{g\cdot j}H$ for all $g\in G$ and all $1\leq j\leq n$. The action is faithful and transitive. We have ${\rm Stab}(x_{j}H)=x_{j}Hx_{j}^{-1}$. Since the action is faithful, it follows that $\bigcap\limits_{1\leq j\leq n}(x_{j}Hx_{j}^{-1})=\bigcap\limits_{1\leq j\leq n}({\rm Stab}(x_{j}H))\ =\{1\}.$
Let $K_{G}/K$ be finite Galois extension with Galois group isomorphic to $G$. We identify it with $G$. Let $L$ be the subfield of $K_{G}$ fixed by $H$. The Galois closure of $L/K$ in $K_{G}$ is the subfield of $K_{G}$ corresponding to the intersection of the conjugates of $H$ in $G$ and that intersection is trivial because of argument in previous paragraph. Hence $K_{G}$ is the Galois closure of $L/K$. Let $f$ over $K$ be the minimal (hence irreducible) polynomial for a primitive element of $L/K$. We can identify the $n$ roots of $f$ with the $n$ cosets in $G/H$. This is the required polynomial $f$. ∎
@@
As a corollary, we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$. This corollary also holds for $G^{n}$ in place of $G$ for any $n\in\mathbb{N}$ as well as for $K^{\prime}$.
@@
Now for $i\neq j$, $E_{i}\cap E_{j}$ corresponds to subgroup generated by $N_{i}$ and $N_{j}$ which is $G^{n}$. Hence $E_{i}\cap E_{j}=K$. Suppose for some $i\neq j$ we have $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}$. Since $E_{i}\cap E_{j}=K$, we get $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}=K$. Now suppose that all $E_{i}\cap K^{\prime}$ are distinct. Since we have $n$ intermediate fields of $K^{\prime}/K$, $E_{i}\cap K^{\prime}=K$ for some $i$. In either case we get an $i$ such that $E_{i}\cap K^{\prime}=K$. Hence ${\rm Gal}(E_{i}K^{\prime}/K^{\prime})\cong{\rm Gal}(E_{i}/K)\cong G$. This realises $G$ as a Galois group over $K^{\prime}$. By replacing $G$ with $G^{m}$ for any $m\in\mathbb{N}$ in the above argument, we can realise $G^{m}$ over $K^{\prime}$ for any $m$.
For the proof of the corollary, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$. We observe that $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$. Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$ with $G$ as Galois group of each $E_{i}/K$. ∎
@@
Now suppose $r>1$. In solutions of Exercises 3 and 4 in [7], a solvable group $G\subseteq{\mathfrak{S}}_{n}$ is constructed with the properties that its action is transitive on $n$ points, and a point stabiliser fixes precisely $r$ points. The construction is as follows: We divide the $n$ points into $n/r=s$ packets of size $r$. Let $G$ be the group of permutations on these $n$ points generated by independent cyclic permutations on each packet, together with a cyclic permutation on the overall set of packets. Hence $G$ is transitive. This construction of $G$ has the explicit description of a semidirect product of an $s$-fold direct product of cyclic groups ${\mathbb{Z}}/r{\mathbb{Z}}$ and a cyclic group ${\mathbb{Z}}/s{\mathbb{Z}}$. A semidirect product group law on $G$ is given by
$((a_{1},\dots,a_{s}),b)\cdot((c_{1},\dots,c_{s}),d)=((a_{1},\dots,a_{s})+(b\cdot(c_{1},\dots,c_{s})),b+d),$
where $b\cdot(c_{1},\dots,c_{s})=(c_{b+1},\dots,c_{s},c_{1},\dots,c_{b})$.
Thus, $G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}},$ where each ${\mathbb{Z}}/r{\mathbb{Z}}$ permutes points in a packet and action of ${\mathbb{Z}}/s{\mathbb{Z}}$ permutes $s\ \text{copies of}\ {\mathbb{Z}}/r{\mathbb{Z}}.$
@@
Since direct product of solvable groups is solvable, direct products $G^{i}$ for $i\in\mathbb{N}$ are solvable. By Shafarevich’s theorem ([9]), $G^{i}$ for $i\in\mathbb{N}$ are realisable as Galois groups over $\mathbb{Q}$. Hence by Lemma 3.1.4, $G$ is realisable as Galois group over number field $K$. By Lemma 3.1.3, there exists an irreducible polynomial $f$ over $K$ of degree $n$ and a labelling of the roots so that the Galois group of $f$, viewed as a group permuting roots of $f$, is precisely $G$. This $f$ satisfies $r_{K}(f)=r$.
@@ 3.2. A simple lemma about $s_{K}(L)$
We begin this subsection by giving an alternate proof for last equality in (3) of Thm.2.1.1, which is stated in [8] and proved in first Proposition in unpublished note of Perlis [7]. The equality states: $r_{K}(f)=[N_{G}(H):H].$
@@
Let $L_{1},L_{2},\dots,L_{s_{K}(L)}$ be the distinct fields as in above Lemma 3.2.1. Hence we have $\tilde{L}=L_{1}L_{2}\dots L_{s_{K}(L)}$, that is the Galois closure of $L/K$ is compositum of distinct fields isomorphic to $L$ over $K$.
@@
Let $K$ be a number field. Note that ${}^{n}P_{k}$ and ${}^{n}C_{k}$ are integers with ${}^{n}P_{k}=k!$ ${}^{n}C_{k}$. By Thm. 3.1.1, we get irreducible polynomial over the field with degree ${}^{n}P_{k}$ and cluster size $k!$. The following theorem says that this is also true under some condition for a general perfect field.
By using the above Lemma  3.2.1, we give an alternate proof for Thm. 3 in [5].
@@
Let $K$ be a perfect field. Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$ (For a number field $K$, such $f$ always exists. See $r=1$ case in Proof of Thm. 3.1.1).
For $21\leq k\leq n-$ 2, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$. Let $g$ be the irreducible polynomial over $K$ for a primitive element of $L_{k}$. This polynomial has degree ${}^{n}P_{k}$ and cluster size $k!$.
@@
Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$ be roots of $f$ in $\bar{K}$. Let $L_{k}=K(\alpha_{1},\alpha_{2},\ldots,\alpha_{k})$. We have that degree of $L_{k}/K$ is $n(n-1)\ldots(n-k+1)=$${}^{n}P_{k}$. Since Galois group of $f$ is ${\mathfrak{S}}_{n}$, we have $L^{\prime}$ is isomorphic to $L_{k}$ over $K$ $\iff$ $L^{\prime}=K(\alpha_{i_{1}},\alpha_{i_{2}},\ldots,\alpha_{i_{k}})$ for $k$ roots $\alpha_{i_{j}}:1\leq j\leq k$ of $f$. By Lemma 3.2.1, $s_{K}(L_{k})$ is number of distinct fields inside $\bar{K}$ isomorphic to $L_{k}$ over $K$ which is precisely the number of ways of choosing $k$ roots from $n$ roots which is ${}^{n}C_{k}$. By Thm. 2.1.1 (2), $r_{K}(L_{k})=k!$. ∎